OpenStudy (anonymous):

Helpful steps please, Find the radius of a circle in which the central angle of pi/6 radian determines a sector area of 43 sq meters. Round to the nearest hundredth.

OpenStudy (freckles):

\[\text{ area of sector with central angle theta}=\frac{1}{2}r^2 \theta\]

OpenStudy (anonymous):

So then would the set up be 43 = 1/2 r^2 pi/6

OpenStudy (istim):

Yes. it would be 43 = (1/2) * (r^2)(pi/6) Solving for r, first step, divide by 1/2 or multiply equation by 2: 86 = (r^2)(pi/6) Divide by pi/6 (516/pi) = r^2 Take the square root.. r = sqrt(516/pi) Therefore, radius is sqrt(516/pi)

OpenStudy (anonymous):

oh I just wrote that out myself I got 12. 82 m